OpenStudy (anonymous):

Please help! 4cos (theta) + 1= 2cos (theta)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@robtobey

hero (hero):

\(4\cos(\theta) + 1 = 2\cos(\theta)\) right?

OpenStudy (anonymous):

yes!

hero (hero):

For the moment, let \(x = \cos(\theta)\)

hero (hero):

Now you have \(4x + 1 = 2x\)

hero (hero):

Solve for x

OpenStudy (anonymous):

2x=-1?

hero (hero):

To finish solving for x, divide both sides by 2 right?

OpenStudy (anonymous):

ohh, yes!

OpenStudy (anonymous):

you would get x= -1/2

hero (hero):

Also, I'm going to have to ask you to change your username to something with a more positive connotation.

OpenStudy (anonymous):

lol, will do

hero (hero):

Which means \(\cos(\theta) = -\dfrac{1}{2}\)

hero (hero):

Now take inverse cosine of both sides to isolate \(\theta\)

OpenStudy (anonymous):

so the inverse cosine would be 120 degrees?

OpenStudy (anonymous):

and 240 degrees

OpenStudy (anonymous):

which would be 2pi/3 and 4pi/3?

hero (hero):

2pi/3 is correct

hero (hero):

From where did you get 4pi/3 ?

OpenStudy (anonymous):

from 240 degrees

OpenStudy (anonymous):

i was suppose to find 0< theta< 2pi. that means the whole unit circle right? I forgot to write that in the original question D:

OpenStudy (anonymous):

thank you so much